i need help on this word problem.
amy is trying to eat a healthier diet. she knows that a carrot has 13 more calories than a stalk of celery and that 5 carrots and 10 celery stalks together have only 170 calories. help her find the calories in each food.
thank you : )

Answer by TimothyLamb(4379)

c = cal/celery
k = cal/carrot
---
k = c + 13
5k + 10c = 170
---
5k + 10c = 170
5(c + 13) + 10c = 170
5c + 65 + 10c = 170
15c = 105
---
c = celery = 7 cal
k = carrot = 20 cal
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php